1. Explain what it means for an energy source to be renewable or non-renewable.
  2. A camera weighing 20 N falls from a small drone hovering 35 m overhead and enters free fall. What is the gravitational potential energy change of the camera from the drone to the ground if you take a reference point of (a) the ground being zero gravitational potential energy? (b) The drone being zero gravitational potential energy? What is the gravitational potential energy of the camera (c) before it falls from the drone and (d) after the camera lands on the ground if the reference point of zero gravitational potential energy is taken to be a second person looking out of a building 40 m from the ground?
  3. A sled of mass 80 kg starts from rest and slides down a 20° incline 80 m long. It then travels for 25 m horizontally before starting back up an 9° incline. It travels 80 m along this incline before coming to rest. What is the magnitude of the net work done on the sled by friction?
1
Expert's answer
2021-07-05T08:44:45-0400

1) Renewable resources are an energy source that cannot be depleted and are able to supply a continuous source of clean energy.

2) a)


"\Delta U=(20)(35)=700\ J"

b)


"\Delta U=0\ J"

c)


"\Delta U=(20)(35)-(20)(40)=-100\ J"

d)


"\Delta U=(20)(0)-(20)(40)=-800\ J"

3)


"W=mg(H-h)\\
W=(80)(9.8)(80\sin{20}-80\sin{9})=11600\ J"